# The spectrum of Dirichlet-to-Neumann maps for radial conductivities

### v1

@@ 3.1. Weyl-Titchmarsh functions and DtN maps
In the presence of radial symmetry, it is possible to express $\Lambda^{\mathrm{p}}_{V}$ in terms of the Dirichlet-to-Neumann map of a one-dimensional Schrödinger operator on the half-line:
(3.1) $\begin{cases}-v_{z}^{\prime\prime}+Qv_{z}=-z^{2}v_{z},\qquad\text{ in }\mathbb{R}_{+},\\ v_{z}\in L^{2}(\mathbb{R}_{+}),\end{cases}$
When the potential $Q$ belongs to $L^{1}(\mathbb{R}_{+})$, it is known that there exists $\beta_{Q}>0$ such that, when $z\in\mathbb{C}_{+}\setminus[0,\beta_{Q}]$, problem (3.1) has a unique solution. One then defines the Weyl-Titchmarsh function of $Q$, or Dirichlet-to-Neumann map, as:
(3.2) $m_{Q}(-z^{2}):=\frac{v^{\prime}_{z}(0)}{v_{z}(0)}.$
@@
The connection of $m_{Q}$ with radial Dirichlet-to-Neumann maps is as follows. Given a potential $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p>d/2$ and $Y_{k}\in\mathfrak{H}^{d}_{k}$ then the unique solution $u\in H^{1}(\mathbb{B}^{d})$ of
(3.3) $\begin{cases}(-\Delta+V)u&=0,\quad\text{ in }\mathbb{B}^{d},\\ \hfill u|_{\mathbb{S}^{d-1}}&=Y_{k},\end{cases}$
@@
[…]
where $v_{k+\nu_{d}}\in L^{2}(\mathbb{R}_{+})$ solves the boundary value problem (3.1) for all $k\in\mathbb{N}_{0}$ with potential $Q=Q_{V}$ given by any of the following equivalent identities. If $V=q_{V}(|\cdot|)$ then
@@
Moreover
(3.6) $\lambda_{k}[V]=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-\nu_{d}\qquad\forall k\in\mathbb{N}_{0}.$
@@
$\sigma_{k}[V^{\mathrm{B}}]=\lambda_{k}[V]-k,\qquad\forall k\in\mathbb{N}_{0},\;k>k_{V},$
where
(3.7) $V^{\mathrm{B}}(x):=|x|^{-2}A_{Q_{V}}(-\log|x|),\qquad x\in\mathbb{B}^{d}\setminus\{0\},$
This means that the $A$-amplitude of $Q_{V}$ determines the Born approximation of the potential $V$.
@@
The important formula (3.7) which connects the Born approximation $V^{\mathrm{B}}$ to the $A$-amplitude of the potential $Q_{V}$ will allow us to extract precise information on the singularities of $V^{\mathrm{B}}$ at 0 from certain asymptotic properties of $A_{Q_{V}}$.
@@ 3.2. Asymptotics of $A$-amplitudes
$D(H_{Q})=\{u\in L^{2}(\mathbb{R}_{+})\,:\,u\in\mathcal{A}\mathcal{C}_{\mathrm{loc}}(\mathbb{R}_{+}),\,H_{Q}u\in L^{2}(\mathbb{R}_{+}),\,u(0)=0\}$
@@
These solutions are known as Jost solutions (they are usually defined using the variable $z^{\prime}=iz$), and $F(z):=\psi(0,z)$ is the Jost function associated to $Q$. Also, both functions $F$ and $\psi(t,\cdot)$ are analytic in $\mathbb{C}_{+}$. Notice that the zeros of $F$ at $\mathbb{C}_{+}$ occur at the points $z$ such that $z^{2}$ is a Dirichlet eigenvalue of $H_{Q}$.
One can show that $H_{Q}$ has no eigenvalues in $0,\infty)$, and that it always has a finite number of negative eigenvalues $\kappa_{j}^{2}$, where $j=1,\dots,J$, with $J\in\mathbb{N}_{0}$ and $0<\kappa_{J}<\dots<\kappa_{1}$. Recall that the Jost function associated to $Q$ satisfies $F(\kappa_{j})=0$ for all $j\in J$. It will also be relevant if the Jost function vanishes at $z=0$. In fact, as in [RS00], we will say that $H_{Q}$ has a zero resonance iff $F(0)=0$.
@@
(3.11) $\left\lVert f\right\rVert_{L^{1}_{\ell}(\mathbb{R}_{+})}:=\int_{0}^{\infty}|f(t)|(1+t)^{\ell}\,dt.$
@@ 3.3. Singularities of the Born approximation
We will now describe how the results of the previous section translate into results on the Born approximation for Schrödinger operator Dirichlet-to-Neumann maps. The following Lemma shows, in particular, that as soon as $p>d/2$ and $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ we have that $Q_{V}$ satisfies (3.8).
@@
Let $1/p+1/q=1$. Then
$\left\lVert Q_{V}\right\rVert_{L^{1}_{\alpha}(\mathbb{R}_{+})}=\int_{\mathbb{R}_{+}}|Q_{V}(t)|(1+t)^{\alpha}\,dt$ $=\int_{\mathbb{B}^{d}}|V(x)|(1+|\log|x||)^{\alpha}|x|^{2-d}\,dx$
$\leq\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}\left(\int_{\mathbb{B}^{d}}(1+|\log|x||)^{q\alpha}|x|^{q(2-d)}\,dx\right)^{1/q},$
@@
$V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
@@
In general, each constant $a_{j}$ is explicitly related to certain values of the eigenfunction with eigenvalue $\kappa_{j}$, see [RS00].
@@
Let $d\geq 2$, $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$. Then
(3.15) $\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}(x)||x|^{2k}\,dx<\infty,\quad\text{and}\quad\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{p}}_{k}[V]-k,\qquad\forall k\in\mathcal{N},$
where $\mathcal{N}\subseteq\mathbb{N}_{0}$ is given by:
@@
    If the point spectrum of $H_{Q_{V}}$ is non-empty then $\mathcal{N}=\{k\in\mathbb{N}_{0}:k>\kappa_{1}-\nu_{d}\}$.
@@
      If $d=2$ then $\mathcal{N}=\mathbb{N}_{0}$ provided that $H_{Q_{V}}$ has no zero resonance, and $\mathcal{N}=\mathbb{N}$ otherwise.
@@
$\lambda_{k}[V]-k=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-(k+\nu_{d})\\ =\int_{0}^{\infty}e^{-2kt}A_{Q_{V}}(t)e^{-(d-2)t}\,dt=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}V^{\mathrm{B}}(x)|x|^{2k}\,dx=\sigma_{k}[V^{\mathrm{B}}],$
@@
3.4. The conductivity case
@@
$|w(x)|\leq(|A|+|B|)M(x)\exp\left(\int_{x}^{\infty}t|Q_{V}(s)|ds\right),\qquad M(x):=\int_{x}^{\infty}t(1+t)|Q_{V}(t)|\,dt,$
@@
    If $d>2$, the discrete spectrum of $H_{Q_{V}}$ is contained in $(-\nu_{d}^{2},0)$.
@@
    If $d=2$, the discrete spectrum of $H_{Q_{V}}$ is empty and it has no zero resonance.
@@
Notice that $(u,(H_{Q_{V}}+\nu_{d}^{2})u)_{L^{2}(\mathbb{R}_{+})}$ is always positive in $H^{1}_{0}(\mathbb{R}_{+})$ since the second integral only vanishes for $u=c\phi$ with $c\in\mathbb{C}$, but $\phi(0)=\gamma_{0}(1)>0$, hence $\phi\notin H^{1}_{0}(\mathbb{R}_{+})$. Therefore
@@
By Lemma˜3.8, 0 is not an eigenvalue $H_{Q_{V}}$. By [Mar86, Lemma 3.1.1], for $\kappa\in\mathbb{R}$ the solutions of $v^{\prime\prime}+Q_{V}v=\kappa^{2}v_{z}$ has two independent solutions $v_{\pm}$ that behave as $e^{\pm i\kappa t}$ for $t\to+\infty$, so $\kappa^{2}\in(0,\infty)$ can never be an eigenvalue of $H_{Q_{V}}$. This ensures the absence of embedded eigenvalues in $0,\infty)$ (since $Q_{V}\in L^{1}_{\alpha}(\mathbb{R}_{+})\subset L^{1}(\mathbb{R}_{+})$, one can also apply the classical result of Kneser–Weyl [Kne26, Wey10] –see also Simon [Sim19]– to prove the absence of embedded eigenvalues in $(0,\infty)$). Hence, it follows that the discrete spectrum of $H_{Q_{V}}$ must be contained in $(-\nu_{d}^{2},0)$ if $d>2$, and must be empty if $d=2$.
@@
so if $d=2$ we have $H_{Q_{V}}\phi=0$. Now, $\phi(t)\to\gamma(0)>0$ as $t\to\infty$, so $\phi(t)=\psi(t,0)$ is the Jost solution with $z=0$ associated to $H_{Q_{V}}$ introduced in Section˜3.2. The Jost solution is unique and we have $\psi(0,0)=\phi(0)=\gamma_{0}(1)>0$. This implies that $H_{Q_{V}}$ has no Dirichlet zero resonance. ∎
As the following proposition shows, the previous conditions on the spectrum are enough to prove that $V^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$ always, with some extra regularity properties.
Proposition 3.10.
Let $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$. Let $H_{Q_{V}}$ be the operator defined in (3.9) and $V^{\mathrm{B}}$ as in (3.7). The following assertions hold:
  1. i)
    $V^{\mathrm{B}}\in L^{1}_{\mathrm{rad}}(\mathbb{B}^{d})$ and $\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{p}}_{k}[V]-k$, for all $k\in\mathbb{N}_{0}$.
  2. ii)
    If $d>2$ and the discrete spectrum $(-\kappa_{j}^{2})_{j=1}^{J}$ of $H_{Q_{V}}$ is contained in $\left(-\nu_{d}^{2},0\right)$, then there exist some constants $a_{0}\in\mathbb{R},\,a_{1},\dots,a_{J}\leq 0$, such that
    $V^{\mathrm{B}}(x)=\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}}+V^{\mathrm{B}}_{\mathrm{reg}}(x),$
    where $\left\lVert V^{\mathrm{B}}_{\mathrm{reg}}\right\rVert_{d,\alpha}<\infty$ for all $\alpha\geq 0$ (see (3.14) for the definition of $\left\lVert\cdot\right\rVert_{d,\alpha}$).
  3. iii)
    $d=2$, the discrete spectrum of $H_{Q_{V}}$ is empty and it has no zero resonance. Then $V^{\mathrm{B}}$ satisfies $\left\lVert V^{\mathrm{B}}\right\rVert_{2,\alpha}<\infty$ for all $\alpha\geq 0$.
Proof.
Statement i) follows from Lemma˜3.7. ii) and iii) follow directly from Theorem˜3.4 and Lemma˜3.9 (for $H_{Q_{V}}$ the discrete spectrum and the point spectrum coincide, since the essential spectrum is $0,\infty)$ and 0 is never an eigenvalue). ∎

### v2

@@ 3.1. Weyl-Titchmarsh functions and DtN maps
In the presence of radial symmetry, it is possible to express $\Lambda^{\mathrm{S}}_{V}$ in terms of the Dirichlet-to-Neumann map of a one-dimensional Schrödinger operator on the half-line, as we now show. We start by considering the one-dimensional boundary value problem
(3.1) $\begin{cases}-v_{z}^{\prime\prime}+Qv_{z}=-z^{2}v_{z},\qquad\text{ in }\mathbb{R}_{+},\\ v_{z}\in L^{2}(\mathbb{R}_{+}).\end{cases}$
When the potential $Q$ belongs to $L^{1}(\mathbb{R}_{+})$, it is known that there exists $\beta_{Q}>0$ such that problem (3.1) has a unique solution when $z\in\mathbb{C}_{+}\setminus[0,\beta_{Q}]$. One then defines the Weyl-Titchmarsh function of $Q$ as:
$m_{Q}(-z^{2}):=\frac{v^{\prime}_{z}(0)}{v_{z}(0)}.$
@@
The connection of $m_{Q}$ with the radial Dirichlet-to-Neumann map $\Lambda^{\mathrm{S}}_{V}$ defined in Section˜2.1 is as follows. Given a potential $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p>d/2$ such that zero is not a Dirichlet eigenvalue of $\Delta+V$, and a spherical harmonic $Y_{k}\in\mathfrak{H}_{k,d}$, then the unique solution $u\in H^{1}(\mathbb{B}^{d})$ of
$\begin{cases}(-\Delta+V)u&=0,\quad\text{ in }\mathbb{B}^{d},\\ \hfill u|_{\mathbb{S}^{d-1}}&=Y_{k},\end{cases}$
@@
$u(x)=|x|^{-\nu_{d}}v_{k+\nu_{d}}(-\log|x|)Y_{k}(x/|x|),$
where, for all $k\in\mathbb{N}_{0}$, $v_{k+\nu_{d}}\in L^{2}(\mathbb{R}_{+})$ solves the boundary value problem (3.1) with potential $Q=Q_{V}$ given by any of the following equivalent identities: if $V=q_{V}(|\cdot|)$ then
@@
Then, a direct computation shows that $m_{Q_{V}}$ determines the spectrum of $\Lambda^{\mathrm{S}}_{V}$ by the relation
(3.3) $\lambda^{\mathrm{S}}_{k}[V]=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-\nu_{d}\qquad\forall k\in\mathbb{N}_{0}.$
@@
(3.4) $\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{S}}_{k}[V]-k,\qquad\forall k\in\mathbb{N}_{0},\;k>k_{V},$
where $V^{\mathrm{B}}$ is defined in terms of the $A$-amplitude as
(3.5) $V^{\mathrm{B}}(x):=|x|^{-2}A_{Q_{V}}(-\log|x|),\qquad x\in\mathbb{B}^{d}\setminus\{0\}.$
In addition, it is also proved in [DMMN24, Theorem 1] that $|\cdot|^{2k}V^{\mathrm{B}}\in L^{1}(\mathbb{B}^{d})$.
The main objective in this section is to refine these results when $V$ arises from a conductivity, in order to prove Proposition˜2.1. The important formula (3.5) which connects the Born approximation $V^{\mathrm{B}}$ to the $A$-amplitude of the potential $Q_{V}$ will allow us to extract precise information on the singularities of $V^{\mathrm{B}}$ at 0 from certain asymptotic properties of $A_{Q_{V}}$.
@@ 3.2. Asymptotics of $A$-amplitudes
$D(H_{Q})=\{u\in L^{2}(\mathbb{R}_{+})\,:\,u\in\mathcal{A}\mathcal{C}_{\mathrm{loc}}(\mathbb{R}_{+}),\,H_{Q}u\in L^{2}(\mathbb{R}_{+}),\,u(0)=0\},$
@@
These solutions are known as Jost solutions (they are usually defined using the variable $z^{\prime}=iz$), and $F(z):=\psi(0,z)$ is the Jost function associated to $Q$. In addition, both functions $F$ and $\psi(t,\cdot)$ are analytic in $\mathbb{C}_{+}$. Notice that the zeros of $F$ at $\mathbb{C}_{+}$ occur at the points $z$ such that $z^{2}$ is a Dirichlet eigenvalue of $H_{Q}$.
We will later show that, under suitable assumptions on $Q$, the operator $H_{Q}$ has no eigenvalues in $0,\infty)$. In addition, it always has a finite number of negative eigenvalues $\kappa_{j}^{2}$, where $j=1,\dots,J$, with $J\in\mathbb{N}_{0}$ and $0<\kappa_{J}<\dots<\kappa_{1}$ (see [RS00]). Recall that the Jost function associated to $Q$ satisfies $F(\kappa_{j})=0$ for all $j\in J$. It will also be relevant if the Jost function vanishes at $z=0$. In fact, as in [RS00], we will say that $H_{Q}$ has a zero resonance iff $F(0)=0$.
@@
$\left\lVert f\right\rVert_{L^{1}_{\ell}(\mathbb{R}_{+})}:=\int_{0}^{\infty}|f(t)|(1+t)^{\ell}\,dt.$
@@ 3.3. Singularities of the Born approximation
We will now describe how the results of the previous section translate into results on the Born approximation for radial Dirichlet-to-Neumann maps $\Lambda^{\mathrm{S}}_{V}$. The following Lemma shows, in particular, that as soon as $p>d/2$ and $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ we have that $Q_{V}$ satisfies (3.6).
@@
Let $1/p+1/q=1$ and $c_{d}:=|\mathbb{S}^{d-1}|^{-1}$. Then
$\left\lVert Q_{V}\right\rVert_{L^{1}_{\alpha}(\mathbb{R}_{+})}=\int_{\mathbb{R}_{+}}|Q_{V}(t)|(1+t)^{\alpha}\,dt$ $=c_{d}\int_{\mathbb{B}^{d}}|V(x)|(1+|\log|x||)^{\alpha}|x|^{2-d}\,dx$
$\leq c_{d}\left\lVert V\right\rVert_{L^{p}(\mathbb{B}^{d})}\left(\int_{\mathbb{B}^{d}}(1+|\log|x||)^{q\alpha}|x|^{q(2-d)}\,dx\right)^{1/q},$
@@
$V^{\mathrm{B}}(x)=V^{\mathrm{B}}_{\mathrm{reg}}(x)+\frac{a_{0}}{|x|^{2}}+\sum_{j=1}^{J}\frac{a_{j}}{|x|^{2\kappa_{j}+2}},$
@@
Let $d\geq 2$, $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$ with $p>d/2$. Then for every $k\in\mathcal{N}$
$\int_{\mathbb{B}^{d}}|V^{\mathrm{B}}(x)||x|^{2k}\,dx<\infty,\quad\text{and}\quad\sigma_{k}[V^{\mathrm{B}}]=\lambda^{\mathrm{S}}_{k}[V]-k,$
where $\mathcal{N}\subseteq\mathbb{N}_{0}$ is the set defined as follows:
@@
    If the point spectrum of $H_{Q_{V}}$ is non-empty, $\mathcal{N}=\{k\in\mathbb{N}_{0}:k>\kappa_{1}-\nu_{d}\}$.
@@
      If $d=2$, $\mathcal{N}=\mathbb{N}_{0}$ provided that $H_{Q_{V}}$ has no zero resonance, and $\mathcal{N}=\mathbb{N}$ otherwise.
@@
$\lambda^{\mathrm{S}}_{k}[V]-k=-m_{Q_{V}}\left(-(k+\nu_{d})^{2}\right)-(k+\nu_{d})\\ =\int_{0}^{\infty}e^{-2kt}A_{Q_{V}}(t)e^{-(d-2)t}\,dt=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{B}^{d}}V^{\mathrm{B}}(x)|x|^{2k}\,dx=\sigma_{k}[V^{\mathrm{B}}],$
@@
3.4. The conductivity case and proof of Proposition˜2.1
@@
$|w(x)|\leq(|A|+|B|)M(x)\exp\left(\int_{x}^{\infty}t|Q_{V}(t)|\,dt\right),\qquad M(x):=\int_{x}^{\infty}t(1+t)|Q_{V}(t)|\,dt,$
@@
    If $d>2$, the point spectrum of $H_{Q_{V}}$ is contained in $(-\nu_{d}^{2},0)$.
@@
    If $d=2$, the point spectrum of $H_{Q_{V}}$ is empty and it has no zero resonance.
@@
Notice that $(u,(H_{Q_{V}}+\nu_{d}^{2})u)_{L^{2}(\mathbb{R}_{+})}$ is always positive in $H^{1}_{0}(\mathbb{R}_{+})$ since the second integral only vanishes for $u=c\phi$ with $c\in\mathbb{C}$, but $\phi(0)=\sqrt{\gamma_{0}(1)}>0$, hence $\phi\notin H^{1}_{0}(\mathbb{R}_{+})$. Therefore
@@
By Lemma˜3.8, 0 is not an eigenvalue $H_{Q_{V}}$. By [Mar86, Lemma 3.1.1], for $\kappa\in\mathbb{R}$ the solutions of $v^{\prime\prime}+Q_{V}v=\kappa^{2}v_{z}$ has two independent solutions $v_{\pm}$ that behave as $e^{\pm i\kappa t}$ for $t\to+\infty$, so $\kappa^{2}\in(0,\infty)$ can never be an eigenvalue of $H_{Q_{V}}$. This ensures the absence of embedded eigenvalues in $0,\infty)$ (since $Q_{V}\in L^{1}_{\alpha}(\mathbb{R}_{+})\subset L^{1}(\mathbb{R}_{+})$, one can also apply the classical result of Kneser–Weyl [Kne26, Wey10] –see also Simon [Sim19]– to prove the absence of embedded eigenvalues in $(0,\infty)$). Hence, it follows that the point spectrum of $H_{Q_{V}}$ must be contained in $(-\nu_{d}^{2},0)$ if $d>2$, and must be empty if $d=2$.
@@
so, if $d=2$, we have $H_{Q_{V}}\phi=0$. Since $\phi(t)\to\sqrt{\gamma(0)}>0$ as $t\to\infty$, then
$\psi(t,0)=\frac{1}{\sqrt{\gamma(0)}}\phi(t),$
is the $z=0$ Jost solution associated to $H_{Q_{V}}$ that was introduced in (3.8). The Jost solution is unique and we have $\sqrt{\gamma(0)}\psi(0,0)=\phi(0)=\sqrt{\gamma_{0}(1)}>0$. This implies that $\psi(0,0)>0$ and that $H_{Q_{V}}$ has no Dirichlet zero resonance. ∎
We can now prove Proposition˜2.1.
Proof of Proposition˜2.1.
If $V$ is given by (2.2), then $V\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, since $\gamma(x)>c>0$ in $\mathbb{B}^{d}$. From Theorem˜3.4, Lemma˜3.9, and Lemma˜3.7 it follows there exists a solution $V^{\mathrm{B}}$ of the moment problem (2.4). This solution is unique as follows from [BCMM22, Equation 1.20], which is reproduced here in (2.10). Assertions i) and ii) are immediate from Theorem˜3.4, Lemma˜3.7, and Lemma˜3.9. Assertion iii) is proved in [DMMN24, Theorem 5]. Finally, iv) is an immediate consequence of i) and ii). ∎